Solve for x: 9^x = 3
I think the answer is x=2






Solve for x: 2^9x-4 = 32

Answer by uma(370)

(1) 9^x = 3
==> (3^2)^x = 3
==> 3^(2x) = 3
==> 2x = 1 [ as bases are the same, equate the exponents]
==> x = 1/2 [ divide by 2 on both the sides]
(2) 2^(9x-4) = 32
==> 2^(9x-4) = 2^5 [ as 32 = 2^5]
==> 9x - 4 = 5 [ as bases are the same, equate the exponents ]
==> 9x = 5 + 4
==> 9x = 9
==> x = 1

good luck